OpenStudy (anonymous):

Number graph ranging from negative two to ten on the x and y axes. A line labeled y equals begin fraction three over two end fraction times x is drawn on the graph. A second line, labeled y equals begin fraction negative one over two end fraction x plus four, is drawn on the graph. What is the solution to the system of equations represented by these two lines? A. (2, 0) B. (2, 3) C. (4, 2) D. (0, 4)

OpenStudy (anonymous):

can someone please help me

OpenStudy (danjs):

y = (3/2)*x and y = (-1/2)*x + 4 right

OpenStudy (anonymous):

right

OpenStudy (danjs):

on the graph, the solution will be where they are both in the same place intersection maybe. at a point (x,y)

OpenStudy (danjs):

for two lines, it could be one solution, 0 , or infinity

OpenStudy (anonymous):

ok so c is the answer

OpenStudy (danjs):

you can solve these things in a number of ways, notice here both are already solved for y = y = (3/2)*x y = (-1/2)*x + 4 so you can set the x parts equal for both, since they both are the same, y (3/2)*x = (-1/2)*x + 4

OpenStudy (danjs):

youc an solve that now for x (4/2) x = 4 x = 2

OpenStudy (danjs):

if x=2, then you can just calculate the y value using any equation , they should give same answers

OpenStudy (danjs):

y = 3/2 * 2 = 3 x=2 and y=3, the ordered pair (x,y) = (2,3)

OpenStudy (anonymous):

ok i said so c